OpenStudy (anonymous):

@sammixboo can I get some help with a couple more questions? They are just like the one we just did.

OpenStudy (anonymous):

-4√x-6=-12 I don't know how to make the long looking square root symbol.

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

I'm not sure how I make the long square root symbol

sammixboo (sammixboo):

So it is \(-4\sqrt{x-6} = -12\)

OpenStudy (anonymous):

Yes!

sammixboo (sammixboo):

OK, so any idea on how to start it? Or no

OpenStudy (anonymous):

I would think you would have to do something with the -4 first

jagr2713 (jagr2713):

\[\huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\]

sammixboo (sammixboo):

Yes, you divide both sides of the equation by -4 \(\dfrac{-4\sqrt{x-6}}{-4} = \dfrac{-12}{-4}\) \(\sqrt{x-6} = \dfrac{-12}{-4}\)

sammixboo (sammixboo):

What is -12/-4

OpenStudy (anonymous):

3

sammixboo (sammixboo):

Right! so \(\sqrt{x-6} = 3\) Now do you know what to do?

OpenStudy (anonymous):

Yes! Let me solve

sammixboo (sammixboo):

OK =)

OpenStudy (anonymous):

Would it be 9?

OpenStudy (anonymous):

Could you help me with one more, and then I will quit bugging you! I really appreciate the help.

sammixboo (sammixboo):

Not quite. Wanna give it another try? Or want me to walk you through it

sammixboo (sammixboo):

Oh, I don't mind the bugging! Hahaha. I like doing Square Roots honestly

OpenStudy (anonymous):

Could you walk me through it?

sammixboo (sammixboo):

Sure! So we have \(\sqrt{x-6} = 3\) We now have to square both sides of the equation, remember? After we do that, we have \(x-6 = 3^2\)

sammixboo (sammixboo):

Now solve the exponent, \(3^2\), what is it?

OpenStudy (anonymous):

9

sammixboo (sammixboo):

Correct, so \(x-6 = 9\) Do you know what to do now? =3

OpenStudy (anonymous):

Would it be 15?

sammixboo (sammixboo):

Right!

OpenStudy (anonymous):

Awesome! Thank you so much. Could you help me with one more like that?

sammixboo (sammixboo):

If you still need help, make sure to post it on another post

sammixboo (sammixboo):

Sure!

OpenStudy (anonymous):

So I need to create a new question?

sammixboo (sammixboo):

Yuppers!

OpenStudy (anonymous):

Okay :)